OpenStudy (anonymous):

i have a question that is find a comin denominator it is 567/432.

OpenStudy (lgbasallote):

"common" denominator....i dont think we can find anything "common" in one subject only right?

OpenStudy (lgbasallote):

"common" grammar-wise is comparison...to compare, one needs 2 objects

OpenStudy (pfenn1):

Do you mean that you are supposed to simplify 567/432?

OpenStudy (pfenn1):

Factoring 567 = 3 * 3 * 3 * 3 * 7 432 = 2 * 2 * 2 * 3 * 3 * 3

OpenStudy (pfenn1):

\[\left( 567 \over 432 \right)=\left( (3^4 (7) \over2^3 3^3 \right)=\left( 3(7) \over 2^3 \right)=\left( 21\over 8 \right)\]